# Quantum field theory on curved manifolds

### v1

@@ 4 The Unruh effect and Hawking radiation on manifolds
where $x=a^{-1},t=0$ is the place for defining the simplest local inertial space. If one tries to define a local inertial system for any other point, one has to consider a distorted coordinate system due to Lorentz transformation. [^14] One can see more explanations in Ref.[9], in which figures for the moving frame can be found. As is immediately apparent from the metric, the field equation does not yield the local Stokes phenomenon as was the case with the Schwinger effect[12]. This indicates that there may be a fundamental error in the way of setting the problem. So far, many researchers have truncated locality here and moved on to find a solution in a global space[2, 3], but here we shall stick to the locality of the issue.
Since the metric (covariant derivatives) does not explain the local Stokes phenomenon, the only way left for us is to use the vierbein. The problem is that even if the vierbein is used for the calculation, the same trivial result is obtained once the field equations are written down. Therefore, the vierbein must be used without going through the field equations. Now consider what an accelerating observer would see if the observer looked directly at the vacuum (vacuum solutions) defined in the observer’s local inertial space. The vacuum is observed here by the particle itself, which is ejected from “the vacuum”. Since such particles have no momentum in their unique frame, we can neglect the $x$-dependent component of the vacuum solution. Considering $dt=\cosh\left(a\tau\right)d\tau$ from Eq.(4.8), we have for the time-dependent part;
@@
which have the required Stokes phenomenon. See ref.[12, 14] if the reader is interested in the details about the Stokes phenomenon. One thing that should be noted is that the result does not meet the conventional (global) calculation by a factor of two. This point has to be explained.
@@
To make the point clear, we consider the entanglements that appear in conventional calculations[3]. In our calculation, a difference of the factor of 2 appears in the non-perturbative factors emerging from the Stokes phenomenon. With entanglement, the probability ($P_{1}$ for a particle production) is squared because there are two particles produced. This gives $P_{entangled}=P_{1}^{2}$ in our calculation. In this way, our calculations will give the same results as the conventional calculations if the entanglement is included by hand. In short, if we “assume” entanglement, the two calculations coincide. So what happens if we use our calculations for Hawking radiation? Conventional calculations of the Unruh effect treat infinite inertial space as real, so the existence of such spaces on curved manifolds cannot be assumed. On the other hand, our calculations are based on the definition of manifolds, so the same calculations can be performed in the vicinity of the black hole horizon. When a pair of particles is created in the vicinity of the horizon, one particle inside the horizon can have negative energy (when it is seen from the outside), and the other in the outside can have positive energy. Again, as there are two particles produced, the total production probability is squared to obtain results consistent with Hawking radiation[12, 13]. Thus, if all calculations are performed faithfully to the definition of the manifold, strong doubts will arise about the entanglement of the Unruh effect. [^15] We are not claiming that our analysis is able to provide a proof that there is no entanglement in the global calculations. It is an indisputable fact that entanglement appears in the global calculation.
@@
In the following, we will consider the Unruh effect when the acceleration changes, which is similar to the slowly changing electric field in the Schwinger effect. We are aware of papers claiming that the Unruh effect can be solved for various accelerating motions of the Unruh-DeWitt detector by adjusting the classical orbit of the detector. Again, our analysis cannot provide a proof that these results are wrong, but we believe we are giving a better alternative. Simply because the general calculations of the Unruh-DeWitt detector are somewhat different from the Unruh effect itself, we have treated them as different. See Ref.[14] for more details about what the local calculations of the Unruh-DeWitt detector look like if it is defined on a manifold, where we found the same factor of two discrepancy between the conventional global analysis and our local analysis. Note also that the Unruh-DeWitt detector requires an explicit interaction with the detector to be included, which is unlikely to be applicable directly to Hawking radiation.
@@ 4.1 When the acceleration rate is slowly time-dependent
For later calculations, let us first derive the Rindler coordinate for the case of constant acceleration using a somewhat roundabout approach. For simplicity, we restrict the motion to the x-axis direction. If the acceleration seen by the inertial system is $a(t)$ and the acceleration seen by an observer moving at the speed of $v(t)$ with respect to the inertial system is $a^{\prime}(t^{\prime})$, then the following relationship holds.
@@
Since the function $T$ of $T^{\prime}$ (i.e, $T(T^{\prime})$) is periodic in the direction of the imaginary axis of $T^{\prime}$, we can expect any vacuum function $F(T)$ described by the observer’s time $T^{\prime}$ to be periodic in the observer’s complex time. Intuitively, this suggests that what the accelerating observer sees in the vacuum is expected to be thermal. Here a simple question would arise. The periodic function for imaginary $T^{\prime}$ seen here is merely a parameterization of the elliptic function so that it takes an infinite limit for one of its double periodicity. The simple answer is that this may be a very special case due to the assumption that the acceleration is a constant. Let us see this point in more detail. In the above calculation, we simply had
@@
which gives an elliptic integral after Legendre’s transformation and thus $T$ as the function of $T^{\prime}$ is described by an elliptic function. Obviously, if we consider $\beta\neq 0$, the periodicity of the function $T(T^{\prime})$ is not a simple imaginary. Such “coordinate systems” do not have the special properties of the Rindler coordinates. Therefore, Unlike the conventional Unruh effect for a constant acceleration, the vierbeins cannot be moved to the same form by the Lorentz transformation. This situation is quite similar to the case which appeared when we have considered the weakly time-dependent electric field in the Schwinger effect. Note that our local analysis only considers slices of the local elliptic function at the real axis and do not extrapolate it to infinity. The most familiar example of the elliptic function is probably the motion of a pendulum. Indeed, if we set
@@
for an “oscillation”, [^16] This “oscillation” is not defined for the observer’s time $t^{\prime}$. we can see that the approximate solution at $t=0$ is obtained for the Unruh effect in the similar way as a motion of a pendulum. [^17] Jacobi functions are complex-valued functions of a complex variable $z$ and a parameter $m(=k^{2})$. Using the elliptic integral of the first kind $K(m)$, the Jacobi $sn$ function has two periods $4K(m)$ and $2K(1-m)i$. In the present case we have $m<0$, while for an pendulum it becomes $m>0$. We comment on the case of solving it at other times. If we consider the Unruh effect at $t=t_{0}$, the inertial condition is now $v(t_{0})=0$. We thus have
@@ 5 Conclusions and Discussions
In this paper, the relationship between the Schwinger and Unruh effects has been discussed on the basis of their similarities as theories on manifolds. The two phenomena, which at first sight appear to be the same, turn out to be caused by completely different sources when one looks at the local structure of the manifold. As we expect negative reactions to the idea of defining the vacuum in the tangent space, this part of the article was explained in particular detail. In this paper, the case of a gradual change in curvature has been considered as an example where the correct answer can only be obtained when the vacuum is defined on the tangent space. In our local analysis, the entanglement of the Unruh effect appears to be an apparent one due to the extrapolation of the coordinate system, but this will require further multifaceted verification. The structure of the elliptic function for the complex time, manifested in the Unruh effect, is commonly found in curved spacetime. Further analysis is needed to determine what implication can be gained when such a structure is incorporated into a manifold.
In physics, it is well known that the connection of the electromagnetic $U(1)$ and the metric can be literally unified into the Kaluza-Klein theory. Although in our discussion of the Unruh effect it might have seemed that the Stokes phenomenon was not able to be derived from the metric, it is clear from the Kaluza-Klein theory that it can appear from the metric once they are actually embedded in the Kaluza-Klein theory. On the other hand, concerning the Unruh effect, the situation remains the same in Kaluza-Klein theory, as the Unruh effect is still hidden in analyses using covariant derivatives. The underlying cause of the Unruh effect is that the inertial system is attached in a twisted manner, which cannot be detected by the covariant derivatives (at least locally).
@@
We hope that the local analyses presented in this paper will help people understand the physics of the quantum field theory on curved manifolds.

### v5

@@ 4 The Unruh effect and Hawking radiation on manifolds
one will not be able to find the required Stokes lines. As is immediately apparent from the metric (and the equations), the field equation does not yield the local Stokes phenomenon as was the case with the Schwinger effect[15]. This indicates that there may be a fundamental error in the way of setting the problem. Here, we shall stick to the locality of the issue.
Since the metric (covariant derivatives) does not explain the local Stokes phenomenon, the only way left for us is to use the vierbein. The problem is that even if the vierbein is used for the calculation, the same trivial result is obtained once the field equations are written down, as the Lorentz frame and the inertial frame are not distinguishable by the metric. Therefore, the vierbein must be used without going through the field equations. Now consider what an accelerating observer would see if the observer looked directly at the vacuum (vacuum solutions) defined in the observer’s local inertial space. The vacuum is observed here by the particle itself, which is ejected from “the vacuum”. [^17] Alternatively, one can introduce the Unruh-DeWitt detector as an observer. The problem has been solved in Ref.[17] by paying serious attention to the locality and the differential geometry. The result is consistent with this paper. Since such particles have no momentum in their unique frame, we can neglect the $x$-dependent component of the vacuum solution. Considering $dt=\cosh\left(a\tau\right)d\tau$ from Eq.(4.8), we have for the time-dependent part;
@@
where $dt$ is needed for the calculation, even if $\omega$ is a constant. The situation is illustrated in Fig.6.
Figure 6: The local vacuum defined in the tangent space is seen by an accelerating observer using the vierbein. The Stokes phenomenon on the Moving frame is illustrated in the bottom. The situation is always the same for the observer if the acceleration is constant. The observer sees always the same Stokes phenomenon on the moving frame, where the subjective vacuum for the observer is defined in the tangent space using the inertial frame of the observer.
@@
We have already analyzed the Stokes phenomenon of the above function in detail in Ref.[15, 17]. As the details of the Stokes phenomenon of the solution are not the issue of this paper, we will present only the results here. See Ref.[15, 17] if the reader is interested in the details about the Stokes phenomenon. One thing that should be noted is that the result does not meet the conventional (global) calculation by a factor of two. This point has to be explained here.
To make the point clear, we consider the entanglements that appear in conventional calculations[6]. In the conventional calculation, the factor of 2 is added when taking a trace on the entanglements in the distant wedge. This is because the particle creation in the right wedge is always accompanied with that in the left wedge as $e^{-\pi\omega/a}b_{\bm{\mathrm{k}}}^{L\dagger}b_{\bm{\mathrm{k}}}^{R\dagger}|0_{L}\rangle\otimes|0_{R}\rangle$, where $b_{\bm{\mathrm{k}}}^{L\dagger}$ and $b_{\bm{\mathrm{k}}}^{R\dagger}$ are the creation operators in the left and the right wedges, respectively. One can see that there is a duplication of the factor, which can be separated as $\left(e^{-\pi\omega/2a}b_{\bm{\mathrm{k}}}^{L\dagger}\right)\left(e^{-\pi\omega/2a}b_{\bm{\mathrm{k}}}^{R\dagger}\right)|0_{L}\rangle\otimes|0_{R}\rangle$. This duplication does not appear if the calculation is local. The entanglement appeared because of the extrapolation beyond the applicable range of the moving frame that defines the observer’s vacuum. Also, to be more rigorous, defining a vacuum in the chart is not a preferred method. In our calculation, we have the factor $e^{-\pi\omega/2a}$ because it is derived from the local Stokes phenomenon. For a local pair creation, we will find $e^{-\pi\omega/a}$ but in our case both particles are observed and there is no trace-out. If we assume the entanglement between the distant wedge, our result can reproduce the usual Unruh effect calculations, but since our local calculation is rigorous as mathematics, there is no need to assume such entanglement with a distant wedge. This is the Bogoliubov coefficient of the state mixing in the vacuum when the vacuum is seen by an accelerating observer. Let us be more precise. Noting that the vierbein connects the inertial frame (vacuum) and the observer, we rewrite $\phi$ in the vacuum as
[…]
where $\omega$ is constant but $dt$ is required to introduce the vierbein in an explicit form. If a vierbein gives $dt=(e(\tau))^{t}_{\tau}d\tau$, mixing of the solutions (when it is seen by an accelerating observer whose proper time is $\tau$) after crossing the Stokes line is written by
[…]
Then, we have
[…]
@@
which suggests that the particle production is seen by $\beta_{\pm}\neq 0$. We call this factor the Bogoliubov coefficient, and this factor should be considered as the Boltzmann factor of the Unruh effect. With entanglement, the probability ($P_{1}$ for a particle production in our calculation) is squared because there are two particles produced but only one particle is observed. This gives $P_{entangled}=P_{1}^{2}$ in our calculation. In this way, our calculations will give the same result as the conventional calculation if the entanglement is included by hand. In short, if we “assume” entanglement, our calculation coincides with the conventional calculation, but obviously our local calculation does not require the entanglement. In our previous paper[17], we have calculated the Boltzmann factor of the Unruh-DeWitt detector and found the same result. Therefore, our local calculations are consistent between the Unruh effect and the Unruh-DeWitt detector, but are not consistent with conventional calculations. The discrepancy is explained by the entanglement. Our local analyses in terms of the differential geometry and the exact WKB allow calculating the probability of single particle observation as the single particle production. Then, our results differ by a factor of two from the usual Unruh Effect and the Unruh-DeWitt detector calculations. So what happens if we apply our calculations to Hawking radiation? Conventional calculations of the Unruh effect treat infinite inertial space as real, so the existence of such spaces cannot be assumed for Hawking radiation. On the other hand, our calculations are always locally defined, so the same calculation of the Unruh effect can be performed in the vicinity of the black hole horizon. When a pair of particles is created in the vicinity of the horizon, one particle inside the horizon can have negative energy (when it is seen from the outside), and the other in the outside can have positive energy. Again, as there are two particles produced but only one particle is observed as radiation, the total production probability for a particle radiation is squared to obtain results consistent with Hawking radiation[15, 16]. In this case, unlike the Unruh effect, we do not have to assume unnatural entanglement at a distance. Thus, if all calculations are performed faithfully to mathematics of the differential geometry, strong doubts will arise about the entanglement of the Unruh effect. [^18] We are not claiming that our analysis is able to provide a proof that there is no entanglement in the conventional global calculations. It is an indisputable fact that entanglement appears in the global calculation, although the mathematical consistency is not reliable because of the extrapolation.
@@
For completeness, we will now consider the Unruh effect as the acceleration changes, which is similar to the slowly changing electric field in the Schwinger effect. Simply because the general calculations of the Unruh-DeWitt detector are somewhat different from the Unruh effect itself, we will treat them as different. See Ref.[17] for more details about what the local calculations of the Unruh-DeWitt detector look like if it is defined locally using the differential geometry and the exact WKB. Note also that the Unruh-DeWitt detector requires an explicit interaction with the detector to be included, which is unlikely to be applicable directly to Hawking radiation.
@@ 4.1 When the acceleration rate is slowly time-dependent
Let us first derive the Rindler coordinate for the case of constant acceleration using a somewhat roundabout approach. For simplicity, we restrict the motion to the x-axis direction. If the acceleration seen by the inertial system is $a(t)$ and the acceleration seen by an observer moving at the speed of $v(t)$ with respect to the inertial system is $a^{\prime}(t^{\prime})$, then the following relationship holds.
@@
Since the function $T$ of $T^{\prime}$ (i.e, $T(T^{\prime})$) is periodic in the direction of the imaginary axis of $T^{\prime}$, we can expect any vacuum function $F(T)$ described by the observer’s time $T^{\prime}$ to be periodic in the observer’s complex time. Here a simple question would arise. The periodic function for imaginary $T^{\prime}$ seen here is merely a parameterization of the elliptic function so that it takes an infinite limit for one of its double periodicity. The simple answer is that this may be a very special case due to the assumption that the acceleration is a constant. Let us see this point in more detail. In the above calculation, we simply had
@@
which gives an elliptic integral after Legendre’s transformation and thus $T$ as the function of $T^{\prime}$ is described by an elliptic function. Obviously, if we consider $\beta\neq 0$, the periodicity of the function $T(T^{\prime})$ is not a simple imaginary. The above “coordinate system” for the slowly varying acceleration do not have the special properties of the Rindler coordinates. More specifically, the situation cannot be always the same for the observer. This means that unlike the conventional Unruh effect for a constant acceleration, the vierbeins cannot be moved to the same form by the Lorentz transformation. This situation is quite similar to the case which appeared when we have considered the weakly time-dependent electric field in the Schwinger effect. Note that our local analysis only considers slices of the local elliptic function at the real axis and do not extrapolate it to infinity. The most familiar example of the elliptic function is probably the motion of a pendulum. Indeed, if we set
@@
for an “oscillation”, [^19] This “oscillation” is not defined for the observer’s time $t^{\prime}$. we can see that the approximate solution at $t=0$ is obtained for the Unruh effect in the similar way as a motion of a pendulum. [^20] Jacobi functions are complex-valued functions of a complex variable $z$ and a parameter $m(=k^{2})$. Using the elliptic integral of the first kind $K(m)$, the Jacobi $sn$ function has two periods $4K(m)$ and $2K(1-m)i$. In the present case we have $m<0$, while for an pendulum it becomes $m>0$. We comment on the case of solving it at other times ($t\neq 0$). If we consider the Unruh effect at $t=t_{0}$, the inertial condition is now $v(t_{0})=0$. We thus have
@@ 5 Conclusions and Discussions
In this paper, the relationship between the Schwinger and Unruh effects has been discussed on the basis of their similarities as theories on manifolds. The two phenomena, which at first sight appear to be the same, turn out to be caused by completely different sources when one looks at the local structure of the manifold. As the idea of defining the vacuum in the local tangent space could be unfamiliar in the analyses of the Schwinger effect and the Unruh effect, this part of the article was explained in particular detail. In this paper, the case of a gradual change in curvature has been considered as an example where the correct answer can only be obtained when the vacuum is defined in the local tangent space. In our local analysis, the entanglement of the Unruh effect appears to be an apparent one due to the extrapolation of the coordinate system. On the other hand, our result agrees with Hawking radiation, where the basic procedures of differential geometry can be used.
We will summarise the reasons why the factor of 2 difference appeared between the conventional calculation of the Unruh effect and ours. As we have described in detail in Sec.4, the direct reason for this difference is the entanglement between distant wedges. This entanglement does not actually exist. The problem arose because of the following problems of the conventional calculation.
  • The moving frame was not moving. This ruins the mathematical consistency of the calculation.
  • The Rindler “vacuum” was defined in the chart, but as we have described in detail in Sec.2, the chart is not a good place for defining a vacuum. One must remember how Lie algebra is introduced in differential geometry.
  • The Rindler chart describes eternal acceleration, but the observer cannot be accelerated eternally in a flat spacetime. Mathematically, there is no problem at this point, as the chart can be chosen in any way. The problem arises when one ignores mathematical consistency and behaves as if the chart is “real”.
  • It must be remembered that the “chart and frame” cannot be used outside the application area. When dealing with an accelerating observer, either frame or chart has to be Lorentz-transformed. Normally the frame is Lorentz-transformed and the frame is called the moving frame. Otherwise the mathematical consistency will be lost outside the application area[12].
  • The most serious problem with defining a vacuum in the chart appears when information about the far future or the far distance (arbitrarily defined by the people who have chosen the chart) is used for the calculation.
  • In the conventional calculation, the quantum states are assumed to be coherent over the whole space-time on the Rindler chart. The information in the far distance is used because of the coherence. The entanglement of the Unruh effect appears in this way. Of course, under normal circumstances, it would be inconceivable to define a state that is coherent to infinity on a chart that is merely a landmark.
  • If the calculation is faithful to differential geometry and no “assumption” is made in physics, the information to be used in physics is normally in the frame and the vierbeins, not in the chart.
Despite its critical importance in this field, differential geometry has not been used in good faith in the analysis of the Unruh effect. The reason is that the procedures of the conventional calculation cannot agree with the basic concepts of differential geometry.
If one attempts to solve the problem faithfully to the concepts of differential geometry, one will find that tools are needed to derive the Unruh effect locally. The fact that such tools were not discovered has postponed the issue of the Unruh effect. The discovery of the local analysis by the exact WKB is a breakthrough in this sense.
Now we would like to rephrase our most important conclusion. “The entanglement between distant wedges of the Unruh effect is an illusion caused by a computational glitch.”
@@
We hope that our local analyses presented in this paper will help people understand more clearly the physics of the quantum field theory on curved manifolds.